# Little By Little: Continual Learning via Self-Activated Sparse Mixture-of-Rank Adaptive Learning

### v1

@@ 4 Experiments
In this paper, we conduct experiments across a diverse set of tasks, including continual learning for both vision-language CLIP models and LLMs, and analyze catastrophic forgetting during fine-tuning. Detailed implementation settings and more experiment results are provided in the Appendix.
Table 2: Comparisons on 5-shot MTIL setting. Following the same protocol as in [85, 81, 48].
    Method Aircraft [50] Caltech101 [25] CIFAR100 [39] DTD [15] EuroSAT [30] Flowers [56] Food [8] MNIST [19] OxfordPet [57] Cars [38] SUN397 [80] Average
CLIP
   Zero-shot [60] 24.3 88.4 68.2 44.6 54.9 71.0 88.5 59.4 89.0 64.7 65.2 65.3
Transfer
   Zero-shot [60] 88.4 68.2 44.6 54.9 71.0 88.5 59.6 89.0 64.7 65.2 69.4
   LwF [43] 72.1 49.2 35.9 44.5 41.1 66.6 50.5 69.0 19.0 51.7 50.0
   LwF-VR [21] 82.2 62.5 40.1 40.1 56.3 80.0 60.9 77.6 40.5 60.8 60.1
   WiSE-FT [78] 77.6 60.0 41.3 39.4 53.0 76.6 58.1 75.5 37.3 58.2 57.7
   ZSCL [92] 84.0 68.1 44.8 46.8 63.6 84.9 61.4 81.4 55.5 62.2 65.3
   MoE-Adapter [85] 87.9 68.2 44.1 48.1 64.7 88.8 69.0 89.1 64.5 65.1 68.9
   RAIL-Primal [81] 88.4 68.2 44.6 54.9 71.0 88.5 59.6 89.0 64.7 65.2 69.4
   CoDyRA [48] 92.4 68.4 45.8 54.5 69.6 87.4 65.2 88.5 64.2 64.5 69.9
   MoRA 92.0 68.8 45.6 53.1 68.6 84.4 64.3 89.8 65.4 64.8 69.7
Average
   LwF [43] 23.5 77.4 43.5 41.7 43.5 52.2 54.6 63.4 68.0 21.3 52.6 49.2
   LwF-VR [21] 24.9 89.1 64.2 53.4 54.3 70.8 79.2 66.5 79.2 44.1 61.6 62.5
   WiSE-FT [78] 32.0 87.7 61.0 55.8 68.1 69.3 76.8 71.5 77.6 42.0 59.3 63.7
   ZSCL [92] 28.2 88.6 66.5 53.5 56.3 73.4 83.1 56.4 82.4 57.5 62.9 64.4
   MoE-Adapter [85] 30.0 89.6 73.9 58.7 69.3 79.3 88.1 76.5 89.1 65.3 65.8 71.4
   RAIL-Primal [81] 32.9 94.5 69.9 58.1 71.8 84.4 88.5 70.4 89.0 66.1 65.7 71.9
   CoDyRA [48] 34.6 95.8 73.9 60.0 77.1 81.3 86.6 75.9 89.9 66.1 65.3 73.3
   MoRA 36.7 95.4 74.9 61.9 77.1 82.6 85.3 76.0 90.5 67.0 65.6 73.9
Last
   LwF [43] 22.1 58.2 17.9 32.1 28.1 66.7 46.0 84.3 64.1 31.5 60.1 46.5
   LwF-VR [21] 22.9 89.8 59.3 57.1 57.6 79.2 78.3 77.7 83.6 60.1 69.8 66.9
   WiSE-FT [78] 30.8 88.9 59.6 60.3 80.9 81.7 77.1 94.9 83.2 62.8 70.0 71.9
   ZSCL [92] 26.8 88.5 63.7 55.7 60.2 82.1 82.6 58.6 85.9 66.7 70.4 67.4
   MoE-Adapter [85] 30.1 89.3 74.9 64.0 82.3 89.4 87.1 89.0 89.1 69.5 72.5 76.1
   RAIL-Primal [81] 32.9 95.1 70.3 63.2 81.5 95.6 88.5 89.7 89.0 72.5 71.0 77.2
   CoDyRA [48] 31.6 95.5 72.8 63.5 85.0 89.7 85.0 94.7 93.2 73.6 73.0 78.0
   MoRA 32.5 95.3 75.3 66.6 87.8 92.6 86.3 96.3 92.6 73.5 73.8 79.3
CL of CLIP We evaluate on two benchmarks using CLIP-ViT/B-16 [60, 33]: MTIL [92, 85], with 11 datasets learned sequentially under a 5-shot split, and X-TAIL [81, 48], with a 16-shot split. Following prior work, we report Transfer (zero-shot on unseen tasks), Last (retention on earlier tasks), and Average (mean of Transfer and Last) to assess continual learning performance.
CL of LLM. We follow prior work [59, 62, 74, 58] to continually fine-tune T5-large [61] on five text-classification benchmarks—AG News, Amazon Reviews, Yelp Reviews, DBpedia, and Yahoo Answers—under three different task orderings. We report the average final accuracy across all tasks after completing the last task, measuring the model’s ability to retain knowledge during CL.
Generalization and forgetting on unseen tasks. To assess effects on pre-trained general knowledge, we fine-tune Llama3 [28] on the CodeAlpaca code-generation dataset [9] and evaluate zero-shot in-domain performance on HumanEval [12], as well as out-of-domain accuracy on a broad selection of MMLU [31] subjects—Formal Logic, Philosophy, World Religions, Economics, Public Relations, STEM, Physics, and Machine Learning. These results reveal how fine-tuning impacts both task-specific adaptation and catastrophic forgetting or improved generalization to unseen domains.
@@ 4.1 Continual Learning of CLIP
Cross-domain task-agnostic incremental learning. Table 1 presents our results on the X-TAIL benchmark. Prior methods [85, 81] rely on zero-shot predictions to preserve pre-trained capabilities and mitigate forgetting, and as a result are capped by the base model’s zero-shot performance. Similarly as done in previous works [92, 48], MoRA aims to continuously refines the pre-trained model, enabling it to surpass this zero-shot ceiling and improve the generalization of the pre-trained model on unseen domains. In practice, continual parameter updates can slightly perturb some zero-shot predictions, our method ultimately improves overall performance on unseen domains by adapting the pre-trained model to new tasks while enhancing its generalization capabilities.
Table 3: Summary of the results on standard CL benchmarks with T5-large model. The averaged accuracy after training on the last task is reported. Experiments on long task sequences in Table 9.
@@
MoRA 77.43 77.47 77.93 77.61
On the Last and Average metrics, MoRA significantly outperforms prior methods. Unlike approaches that require domain ID prediction [85], high dimensional projection layers, or a memory bank of feature embeddings [81], MoRA enhances the pre-trained model’s own representation capability without relying on these external mechanisms. In contrast to the previous state of the art [48], which learns a fixed weight for each rank-$r$ update from each task, MoRA’s novel mixture-of-ranks framework selectively assigns each rank an input-dependent gate. Notably, only a small subset of ranks most relevant to each token are activated, resulting in far more fine-grained utilization and specialization of each ranks.
Multi-domain task incremental learning. We evaluate MoRA in the few-shot MTIL setting (Table 2) under the same protocols as [85, 81, 48]. Consistent with the results observed in the X-TAIL setting, our method demonstrates clear superiority in this scenario.
4.2 Continual Learning of LLM
@@
Table 3 reports results across three task orderings, where MoRA consistently outperforms prior methods and nearly matches the multi-task learning (MTL) upper bound. In contrast to state-of-the-art techniques such as O-LoRA [74] and LB-CL [58], which rely on orthogonality constraints or gradient projections between per-task LoRA adapters—potentially limiting each adapter’s capacity—MoRA requires no additional regularization. By decomposing each rank-$r$ update into its rank-1 components and applying self-activated, sparse gating, MoRA allows every component to specialize on its own input distribution. This fine-grained specialization both prevents interference across tasks and more effectively captures diverse patterns.
4.3 Forgetting and Generalization during Fine-Tuning
@@
Table 4 examines how standard fine-tuning affects both in-domain performance and out-of-domain generalization. We observe two key phenomena: 1. Forgetting of unrelated knowledge. Fine-tuning with a conventional LoRA leads to performance drops on topics completely unrelated to the fine-tuning data—for example, Philosophy, World Religions, and Public Relations—indicating that the model loses pre-trained general knowledge when adapting to a new domain. 2. Improved generalization. Fine-tuning on code generation data also boosts accuracy on tasks requiring logical understanding (Formal Logic, Economics, STEM, Physics, and Machine Learning), suggesting that representations learned during code fine-tuning transfer to related symbolic and quantitative domains.
@@
By contrast, MoRA’s self-routed, sparse mixture-of-ranks framework both preserves unrelated general knowledge and enhances in-domain and out-of-domain generalization performance. By decomposing each low-rank update into specialized rank-1 components and activating only the most relevant subset, MoRA prevents catastrophic overwriting of prior knowledge. At the same time, it achieves strong code-generation performance using only one-third of the activated parameters required by a rank-32 LoRA, and significantly improves the out-of-domain generalization.
Table 4: Fine-tuning Llama3.1-8B on the code generation domain (CodeAlpaca). We report zero-shot in-domain performance on HumanEval (Pass@1) and zero-shot out-of-domain accuracy for selected MMLU subjects—Formal Logic, Philosophy, World Religions, Economics, Public Relations, STEM, Physics, and Machine Learning. The last two columns list trainable parameters (for MoRA: added / activated). The best and second best scores are highlighted in red and blue, respectively.
[…]
Logic Phil. Reli. Econ. Pub. Rel. STEM Phys. ML MMLU
@@
Figure 3: Visualization of MoRA rank activations during Task 1 and Task 2 training. Activations are extracted from the K projection in the attention module (layer 8) of the image encoder. Corresponding image patches are shown below each activation map, with regions relevant to each class marked by orange bounding boxes. Zoom in for details. More visualizations are in Figs. 5 and 6 of Appendix.
4.4 Visualizations on Fine-grained Rank Activations
We visualize rank activations during continual learning in Fig. 3. These visualizations demonstrate that MoRA (1) encourages each rank to specialize in its own input distribution, and (2) reduces cross-task interference to mitigate forgetting.
Each rank specializes in distinct input patterns. In Fig. 6(a), we observed that patches corresponding to the airplane object strongly activate rank 0, whereas background patches (blue sky) predominantly activate rank 11. Similarly as in Fig. 6(c), ranks at index 19, 20, and 29 (highlighted with orange bounding boxes) are activated for image patches of jaguar. Moreover, the complex background elements in Fig. 6(c)—such as tree leaves and shadows—elicit a more diverse activation pattern than the simple blue-sky backgrounds in Fig. 6(a), illustrating MoRA’s capacity to capture varying contextual complexity. Interestingly, we also observe that certain ranks learned on earlier tasks remain active when processing later-task inputs, indicating effective knowledge reuse.
Reduced cross-task interference to mitigate forgetting. By comparing the activation patterns in Fig. 6(a) (after learning Task 1) and Fig. 6(b) (after learning Task 2) on the same input image, we observe that the ranks activations are nearly identical: rank 0 continues to respond to airplane semantics, and rank 11 continues to respond to background blue-sky patches. This consistency demonstrates that our self-activated sparse mixture-of-ranks framework effectively prevents later-task updates from overwriting or intruding upon the knowledge learned for earlier tasks. As a result, inter-task interference is greatly reduced, and catastrophic forgetting is mitigated.
Table 5: Comparisons on different per LoRA/rank mixture calculation strategy
@@ 4.4 Visualizations on Fine-grained Rank Activations
Router@LoRA 62.56 69.45 74.53
Router@rank 60.09 65.97 69.76
Self-activated@rank 60.26 65.94 69.85
     + Sparse activation 62.59 72.28 80.22
     + Adaptive sparse (MoRA) 63.30 72.70 80.90
@@
4.5 Ablation Studies
Per-LoRA/rank mixture weights. Table 5 compares different strategies for computing mixture weights, under a controlled setup with one LoRA per task, standard training loss only, and no additional regularization: (1) Router@LoRA trains a standard router to assign weights to each LoRA adapter. (2) Router@rank trains a fine-grained router for every rank-1 component, which increases training complexity and underperforms Router@LoRA. (3) Self-activated@rank directly infer mixture weights using softmax-normalized activations (Eq. 5), matching or exceeding Router@rank and showing that each rank can infer its own relevance without extra parameters. (4) Sparse activation adds temperature scaling (Eq. 6) and caps the number of activated ranks, yielding substantial gains by sharpening rank specialization. (5) Adaptive sparse activation further prunes low-activation ranks at test time, selecting the most relevant number of ranks to be activated adaptively for each input token, reducing noise from less relevant ranks and achieving the best overall performance.
Rank activation budget. Under our sparse rank-activation framework, we explore how many ranks are necessary by varying the activation budget (Fig. 4(a)). To focus exclusively on this budget parameter, we disable other sparsity enhancements, e.g. rank pruning. After training on 10 tasks (160 ranks total), raising the budget from 2 (1.25%) to 16 active ranks (10%) yields a dramatic boost in final accuracy, and gradually plateaus. This demonstrates that each rank specializes in a distinct input distribution, and a small, sparse subset of ranks suffices to handle diverse input distributions. Moreover, even when the budget is set very high, our self-activating gating remains robust: it still activates only the most relevant ranks for each token while suppressing contributions from less relevant ones.
Temperature setting of $\tau_{\text{MoRA}}$. We use temperature scaling (Eq. (6)) to control mixture sharpness: lower $\tau_{\text{MoRA}}$ yields sharper activations, while higher $\tau_{\text{MoRA}}$ flattens the distribution and engages more ranks. Figure4(b) shows that increasing $\tau_{\text{MoRA}}$ improves transfer to unseen tasks, even surpassing the reported state-of-the-art Transfer in Table 1 when $\tau_{\text{MoRA}}$ is high. A moderate $\tau_{\text{MoRA}}$ around 0.01 balances specificity and generalization, delivering strong Last and Average scores.
Adaptive mixture with rank pruning. In Fig. 4(c), we demonstrate the effects of adaptive sparsity constraints by varying the threshold $\delta$ for rank pruning. Pruning ranks with a relatively small threshold reduces noise and sharpens the mixture, improving both downstream adaptation and out-of-domain generalization, whereas raising $\delta$ above 0.2 begins to remove useful ranks and degrades performance.
@@ 5 Conclusion
We present MoRA, a self-activated sparse mixture-of-ranks framework that decomposes each LoRA update into rank-1 experts and self-infer their sparse mixture weights. By reducing knowledge redundancy and resolving router–expert mismatch during CL, MoRA delivers superior continual-learning performance, significantly reduces forgetting, and enhances generalization.
@@
Limitations and future works. MoRA use a pre-defined temperature to control the sharpness of the mixture weights. While effective, the sparse activation mechanism can be further improved by incorporating awareness of the global data distribution. Learning this temperature or adapting the mixture’s sharpness based on data offers a promising avenue for future research. We also aim to extend the method’s applicability to a broader range of PTMs and applications.
Acknowledgments and Disclosure of Funding
This work was partially supported by the ARC DECRA Fellowship (DE230101591) awarded to D. Gong. H. Lu is affiliated with CSIRO Data61 through a PhD scholarship and acknowledges the support of the Google PhD Fellowship.
Appendix A Additional Details of MoRA
We provide more details of developing the proposed MoRA in this section.
A.1 Model Weights and Updates as Linear Associative Memory Model
Weight matrix as linear associative memory. Large pre-trained models (PTMs) have been shown to inhabit a low “intrinsic-dimension” manifold compared to their full parameter count [41, 1]. Any weight matrix of a pre-trained model $\mathbf{W}_{0}\in\mathbb{R}^{d_{\text{out}}\times d_{\text{in}}}$ can be treated as a linear associative memory model with a key-value structure [37, 5, 6, 54]. Concretely, for any matrix $\mathbf{W}\in\mathbb{R}^{d_{\mathrm{out}}\times d_{\mathrm{in}}}$ and input $\mathbf{x}\in\mathbb{R}^{d_{\mathrm{in}}}$, the product $\mathbf{W}\,\mathbf{x}$ implements a content-addressable memory read over $m$ stored key–value pairs $\{(\mathbf{K}_{i},\mathbf{V}_{i})\}_{i=1}^{m}$, where $m$ denotes the number of key-value memories of a specific weight matrix. Here, ${\mathbf{K}}=\begin{bmatrix}{\mathbf{K}}_{1}^{\top}&\cdots&{\mathbf{K}}_{m}^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{m\times d_{\text{in}}}$ and ${\mathbf{V}}=$$\begin{bmatrix}{\mathbf{V}}_{1}&\cdots&{\mathbf{V}}_{m}\end{bmatrix}\in\mathbb{R}^{d_{\text{out}}\times m}$ with $\mathbf{K}_{i}\in\mathbb{R}^{1\times d_{\text{in}}}$ is the $i$-th row of ${\mathbf{K}}$ and ${\mathbf{V}}_{i}\in\mathbb{R}^{d_{\text{out}}\times 1}$ is the $i$-th column of ${\mathbf{V}}$. Given any input $\mathbf{x}$, the output $\mathbf{y}\in\mathbb{R}^{d_{\text{out}}}$ is computed as
$\mathbf{y}=\mathbf{W}\mathbf{x}\approx\sum\nolimits_{i=1}^{m}\mathbf{V}_{i}(\mathbf{K}_{i}\mathbf{x}).$ (8)
In this formulation, the forward pass through the weight matrix with the input decomposes into two steps: (1) computing response scores or assignment weights via the key vectors $\mathbf{K}_{i}$, and (2) aggregating the corresponding value vectors $\mathbf{V}_{i}$ according to those weights.
Low-rank update as modifications to the linear associative memory. Building on the associative-memory view of weights, we can likewise interpret weight updates during training as key–value insertions into the existing memory. During training, for an arbitrary update $\Delta\mathbf{W}$ on a given $\mathbf{W}_{0}$ as the starting point, the weight update process can be represent as
$\mathbf{W}_{\text{new}}=\mathbf{W}_{0}+\Delta\mathbf{W},$ (9)
Low-Rank Adaptation (LoRA) [32] performs parameter-efficient fine-tuning by adding a rank-$r$ update $\Delta\mathbf{W}=\mathbf{B}\mathbf{A}$ to the pre-trained weight matrix $\mathbf{W}_{0}$, where $\mathbf{B}\in\mathbb{R}^{d_{\text{out}}\times r}$ and $\mathbf{A}\in\mathbb{R}^{r\times d_{in}}$. With the model update based on LoRA, given an input $\mathbf{x}$, the computation of updated weight becomes
Decomposing the low-rank matrices into a collection of its $r$ rank-1 components gives
$\Delta\mathbf{W}\mathbf{x}=\mathbf{B}\mathbf{A}\mathbf{x}=\sum\nolimits_{i=1}^{r}\mathbf{B}_{i}(\mathbf{A}_{i}\mathbf{x}),$ (12)
with $\mathbf{A}=\begin{bmatrix}\mathbf{A}_{1}^{\top}&\cdots&\mathbf{A}_{r}^{\top}\end{bmatrix}^{\top}\in\mathbb{R}^{r\times d_{\text{in}}}$ where $\mathbf{A}_{i}$ is the $i$-th row of $\mathbf{A}$, and $\mathbf{B}=\begin{bmatrix}\mathbf{B}_{1}&\cdots&\mathbf{B}_{r}\end{bmatrix}\in\mathbb{R}^{d_{\text{out}}\times r}$ where $\mathbf{B}_{i}$ is the $i$-th column of $\mathbf{B}$.
Considering the connection between the formulation of LoRA Eq. (1) and associative-memory view of weight update Eq. (10), we can naturally interpret LoRA as maintaining a small key–value store in a low-rank subspace, where the rows of $\mathbf{A}$ serve as keys and the columns of $\mathbf{B}$ serve as their corresponding values. After training the LoRA module on a base weight matrix $\mathbf{W}_{0}$, the modified forward pass in Eq. (11) with Eq. (12) can be seen as: first, the original pre-trained computation $\mathbf{W}_{0}\,\mathbf{x}$ (Eq. (8)) is performed; then, each learned key $\mathbf{A}_{i}$ measures its relevance to the input $\mathbf{x}$ via the product $\mathbf{A}_{i}\mathbf{x}$, and each learned value $\mathbf{B}_{i}$ is scaled by this relevance score and aggregated, in a same way as shown in Eq. (8). In this way, LoRA effectively inserts or updates a small set of key–value memories that augment the pre-trained model’s capabilities.
A.2 Sparse Mixture-of-Rank Adaptation
In the key–value associative memory view of LoRA (Eq. (11) and Eq. (12)), a standard LoRA or its variants [53, 87] apply a dense, fixed update: every rank’s subspace contributes for any kinds of input $\mathbf{x}$. Even in MoE-LoRA, where each LoRA module is treated as a separate expert, all key-value pairs within a LoRA module are used densely as a whole, either fully activated or not activated at all. This leads to redundancy across experts and overlooked the strong potential of the key vectors $\mathbf{A}$ on expressing its relevance to the input.
Rather than activating all key-value memories of each LoRA for all kinds of inputs, where many of which may encode patterns irrelevant to the current input, we introduce a gated, sparse mixture-of-ranks update that selectively activates only the most relevant key–value pairs for each input. Concretely, we associate each rank-1 key–value pair with a scalar weight $w_{i}\in\mathbb{R}$, so that for an input token $\mathbf{x}\in\mathbb{R}^{d_{\mathrm{in}}}$ from task $t$, with $r_{t}$ total ranks, the update becomes:
$\Delta\mathbf{W}^{t}=\sum\nolimits_{i=1}^{r_{t}}w_{i}\mathbf{B}_{i}\mathbf{A}_{i},~{}~{}\text{and}~{}~{}\Delta\mathbf{W}^{t}\mathbf{x}=\sum\nolimits_{i=1}^{r_{t}}w_{i}\mathbf{B}_{i}\mathbf{A}_{i}\mathbf{x},$ (4)
where $w_{i}$ modulates whether and how strongly the corresponding rank contributes to the final updated weight matrix.
As discussed in Sec. 3.4, rather than obtaining mixture weights $w_{i}$’s through learning an additional router [85, 72], considering that $\mathbf{A}_{i}$’s in $\mathbf{A}$ works as keys in the memory model that reflect the significance on the correlation between input and the learned knowledge in each rank, we leverage the key vectors $\mathbf{A}_{i}$themselves to infer each rank’s relevance. Specifically, $w_{i}$’s are obtained through sparsity-inducing scaling with top-k masking and rank pruning of each rank’s activation $\mathbf{A}_{i}\mathbf{x}$, i.e., $[w_{i}]=\texttt{SparseOperator}([\mathbf{A}_{i}\mathbf{x}])$, where $\texttt{SparseOperator}()$ denotes the sparse-inducing operations.
Our fine-grained mixture framework offers two key advantages: (1) It reduces redundancy and enables shared-knowledge reuse: ranks that capture common, cross-task features (such as generic semantics) can be reactivated across multiple tasks, preventing those patterns from being relearned redundantly or interfered; (2) It minimizes routing ambiguity and prevents forgetting: each rank can specialize on more nuanced patterns, yielding an optimized, rank-wise mixture that integrates new knowledge without disrupting existing representations or hindering future updates.
Appendix B Additional Details and Results of Experiments

### v2

@@ 4 Experiments
In this paper, we conduct experiments across a diverse set of tasks, including continual learning for both vision-language CLIP models and LMs, and analyze catastrophic forgetting during fine‐tuning. Detailed implementation settings and more experiment results are provided in Appendix A.1.
CL of CLIP. We evaluate on two benchmarks using CLIP-ViT/B-16 (Radford et al., 2021; Ilharco et al., 2021) on MTIL (Zheng et al., 2023; Yu et al., 2024) and X-TAIL (Xu et al., 2024; Lu et al., 2024), with a range of image classification datasets. Following prior work, we report Transfer (zero-shot on unseen tasks), Last (retention on earlier tasks), and Average (mean accuracy across all datasets of all learning tasks) to assess continual learning performance.
CL of LM. We follow prior works (Qin & Joty, 2021; Razdaibiedina et al., 2023; Wang et al., 2023; Qiao & Mahdavi, 2024) to continually fine-tune T5-large (Raffel et al., 2020) and LLaMA2-7B (Touvron et al., 2023) on five text-classification benchmarks, under three different task orderings. We report the average final accuracy across all tasks after completing the last task.
Generalization and forgetting on unseen tasks. To assess effects on pre-trained general knowledge, we fine-tune Llama3.1-8B (Grattafiori et al., 2024) on the CodeAlpaca code-generation dataset (Chaudhary, 2023) and evaluate zero-shot in-domain performance on HumanEval (Chen et al., 2021), as well as out-of-domain accuracy on a broad selection of MMLU (Hendrycks et al., 2021) subjects.
Table 2: Zero-shot in-domain performance on HumanEval (Pass@1) for the code generation domain and out-of-domain accuracy on selected MMLU subjects (formal logic, philosophy, world religions, economics, public relations, STEM, physics, machine learning) after fine-tuning Llama-3.1-8B on CodeAlpaca. The last two columns report trainable parameters (for MoRA: added / activated).
Method
HumanEval
(Pass@1)
Out-of-Domain (Acc.)
Params
[…]
Logic Phil. Reli. Econ. Pub. Rel. STEM Phys. ML MMLU
@@ 4.1 Continual Learning of CLIP
Table 1 reports X-TAIL results (MTIL results in Appendix Table 11). Prior methods (Yu et al., 2024; Xu et al., 2024) preserve pre-trained capabilities via zero-shot predictions and are therefore capped by the base model’s zero-shot performance. Following recent practice (Zheng et al., 2023; Lu et al., 2024), MoRA continuously trains the pre-trained model so it can surpass that zero-shot ceiling and better generalize to unseen domains, while still preserving existing capabilities. Across both Last and Average metrics, MoRA substantially outperforms prior works. Unlike methods that rely on domain-ID prediction (Yu et al., 2024), high-dimensional projection layers or feature-banks (Xu et al., 2024), MoRA improves the pre-trained representation without external mechanisms; compared with (Lu et al., 2024), which learns fixed weights per rank-$r$ update, MoRA uses input-dependent gates on individual rank-one updates for finer expert specialization and dynamic utilization.
Table 3: Summary of results on standard CL benchmarks with T5-large. We report averaged accuracy after training on the last task across three task orderings.
@@
MoRA 77.4 77.5 77.9 77.6
4.2 Continual Learning of LM
@@
Table 3 reports results across three task orderings: MoRA consistently outperforms prior methods and closely approaches the multi-task learning (MTL) upper bound. Unlike O-LoRA (Wang et al., 2023) and LB-CL (Qiao & Mahdavi, 2024), which rely on orthogonality constraints or gradient projections between per-task LoRA adapters (potentially limiting adapter capacity), MoRA needs no extra regularization. By decomposing each rank-r update into rank-one components and applying self-activated, sparse gating, MoRA lets each component specialize on its own input distribution, reducing interference and more effectively capturing diverse patterns. We further evaluate MoRA on LLaMA2-7B (Touvron et al., 2023) under the same continual-learning setup (Table 9 in the Appendix), MoRA also outperforms O-LoRA by 2.3% averaged over 3 task orders.
4.3 Forgetting and Generalization during Fine‐Tuning
@@
Table 2 examines how standard fine-tuning affects both in-domain performance and out-of-domain generalization. We observe two key phenomena. (1) Forgetting of unrelated knowledge: conventional LoRA fine-tuning causes drops on topics that are unrelated to the fine-tuning data (e.g., Philosophy, World Religions, Public Relations), indicating that pre-trained general knowledge is overwritten during adaptation. (2) Improved generalization: fine-tuning on code data increases accuracy on logically and quantitatively oriented tasks (Formal Logic, Economics, STEM, Physics, Machine Learning), showing that representations learned from code transfer to related symbolic domains.
@@
In contrast, MoRA both preserves unrelated pre-trained knowledge and improves in- and out-of-domain generalization. Decomposing each low-rank update into specialized rank-one components and activating only the most relevant subset prevents catastrophic overwriting of prior knowledge while allowing targeted adaptation. At the same time, MoRA achieves strong code-generation performance using roughly one-third of the activated parameters required by a rank-32 LoRA and substantially improves out-of-domain accuracy.
4.4 Visualizations of Fine-grained Rank Activations
Figure 3 shows rank activations recorded during continual learning. These visualizations illustrate two properties of MoRA: (1) individual ranks specialize on distinct input patterns, and (2) the model substantially reduces cross-task interference, thereby mitigating forgetting. Extended visualizations across more tasks and scenarios appear in Fig. 4 and Fig. 5 in the Appendix.
Each rank specializes in distinct input patterns. In Fig. 3(a), airplane patches strongly activate Rank 0, while blue-sky backgrounds predominantly activate Rank 11. In Fig. 3(c), Ranks 19, 20, and 29 (orange boxes) respond to jaguar patches. The more complex backgrounds in Fig. 3(c) (e.g., leaves, shadows) yield a richer, more distributed pattern than the simple blue sky in Fig. 3(a), highlighting MoRA’s capacity to model contextual complexity. Some ranks learned earlier are also reused on later tasks (Fig. 5, Appendix), indicating transfer of shared semantics.
Reduced cross-task interference and mitigated forgetting. Comparing the same input after Task 1 (Fig. 3(a)) and after Task 2 (Fig. 3(b)) shows almost identical activations (more in Fig. 4, Appendix): Rank 0 still responds to airplane semantics and Rank 11 to blue-sky patches. This stability indicates that our self-activated, sparse mixture-of-ranks prevents later updates from overwriting earlier task representations, reducing inter-task interference and mitigating catastrophic forgetting.
@@ 4.4 Visualizations of Fine-grained Rank Activations
Figure 3: Visualization of MoRA rank activations during Task 1 and Task 2 training. Activations are extracted from the K projection in the attention module (layer 8) of the image encoder. Corresponding image patches are shown below each activation map, with regions relevant to each class marked by orange bounding boxes. Zoom in for details. More visualizations are in Figs. 4 and 5 of the Appendix, demonstrating forgetting mitigation and knowledge reuse.
4.5 Ablation Studies
Routing/activation strategy. Table 4 compares different strategies for computing mixture weights, under a controlled setup with one LoRA per task, standard training loss, and no additional regularization: (1) MoE-LoRA trains a router to assign weights to each LoRA adapter; its coarse-grained design and lack of mechanisms that encourage expert specialization lead to greater interference and increased forgetting compared to MoRA. (2) Non-self-activation trains a finer-grained router for mixing each rank-one component. However, with a much larger number of experts (ranks), the router becomes increasingly prone to confusion, resulting in higher training complexity and degraded performance compared with MoE-LoRA. (3) Self-activation infers weights directly from softmax-normalized activations (Eq. 8), matching or exceeding Non-self-activation and demonstrating that each rank can infer its own relevance without extra trainable router modules. (4) Top-k rank selection restricts updates to the k most relevant ranks, reducing interference and improving accuracy. (5) Sharpness control concentrates activation mass and substantially boosts specialization. (6) Test-time rank thresholding removes low-importance ranks at inference, yielding a slight gain by eliminating noisy contributions.
Table 4: Routing/activation strategy
@@ 4.5 Ablation Studies
MoE-LoRA 62.56 69.45 74.53
Non-self-activation 60.09 65.97 69.76
Self-activation 60.26 65.94 69.85
    w/ top-k rank selection 60.69 66.52 70.62
    w/ sharpness control with $\tau_{\text{MoRA}}$ 62.07 71.15 79.62
    w/ test-time rank thresholding 60.78 66.83 71.08
    MoRA 63.30 72.70 80.90
[…]
@@
Rank activation budget. We probe how many ranks are necessary by varying the activation budget (Fig. 4(a)). To isolate this effect, we disable other sparsity mechanisms (e.g., thresholding). After training on 10 tasks (160 ranks total), increasing the budget from 2 (1.25%) to 16 active ranks (10%) yields a large improvement in final accuracy and then gradually plateaus. This suggests that each rank specializes in distinct input patterns and that a small, sparse subset of ranks is sufficient to cover diverse inputs. Even with a high budget, the self-activating gating remains robust: it still privileges the most relevant ranks per token while suppressing lower-relevance contributions.
Sharpness control of $\tau_{\text{MoRA}}$. We use temperature scaling (Eq. (10)) to control mixture sharpness: lower $\tau_{\text{MoRA}}$ yields sharper activations, while higher $\tau_{\text{MoRA}}$ flattens the distribution and engages more ranks. Fig. 4(b) shows that increasing $\tau_{\text{MoRA}}$ improves transfer to unseen tasks, even surpassing the reported state-of-the-art Transfer in Table 1 when $\tau_{\text{MoRA}}$ is high. A moderate $\tau_{\text{MoRA}}$ around 0.01 balances specificity and generalization, delivering strong Last and Average scores.
Threshold-based rank selection. Figure 4(c) shows the effect of the test-time rank selection threshold $\delta$. Applying a modest threshold removes low-activation ranks and reduces noisy contributions, which improves both downstream adaptation and out-of-domain generalization.
@@ 5 Conclusion
We present MoRA, a self-activated sparse mixture-of-ranks framework that decomposes each LoRA update into rank‐1 experts and self‐infer their sparse mixture weights. By reducing knowledge redundancy and resolving router–expert mismatch during CL, MoRA delivers superior continual‐learning performance, significantly reduces forgetting, and enhances generalization.
@@
Future works. MoRA controls mixture sharpness and sparsity via top-k selection, temperature scaling, and thresholding. While effective, the sparse activation mechanism can be further improved by incorporating awareness of the input data distribution. Learning this temperature or adapting the mixture’s sharpness based on data offers a promising avenue for future research. We also aim to extend the method’s applicability to a broader range of PTMs and applications.
Reproducibility Statement
Further implementation details and dataset descriptions are provided in Sec. 4 and Appendix A.1. Our source code will be publicly released upon paper acceptance.
Appendix A Additional Details and Results of Experiments